You invest $20 at the end of each year for the next 6 years at the annual rate of 9%. What percentage of total interest is due to simple interest?
Answer #1

Total interest = Final amount after 6 years - Initial amount

Total interest = 20*(1.09^6) - 20 = 13.542

Percentage of total interest is due to simple interest = (Total simple interest)*100/Total interest

Percentage of total interest is due to simple interest = (20*6*0.09)*100/13.542 =

Percentage of total interest is due to simple interest = 79.75%